OpenStudy (anonymous):

Find the value of the following definite integral using the limit of a Riemann sum, and express as a common fraction. You should use the formulas for ∑i, ∑i^2, and ∑i^3 given in class. S(Integral from -2 to 1) (x^3-3x)dx

OpenStudy (anonymous):

The actual question is attached

OpenStudy (anonymous):

i attached the actual question

OpenStudy (amistre64):

i see it... so what are the 3 summation formulas?

OpenStudy (anonymous):

let me get them from my notes

OpenStudy (amistre64):

if I recall correctly, the Reiman limits are just taking the area of slices as the thickness of dx goes to zero right?

OpenStudy (anonymous):

yes and i she didnt give us the formulas even though she said she did, let me check the book.

OpenStudy (amistre64):

(f(x+dx) - f(x)) dx; we could also take the "average height" between f(x+dx) and f(x)...

OpenStudy (amistre64):

dx just stands for delta x, any arbitrary number... where dx = (b-a)/n... any of this make sense?

OpenStudy (anonymous):

here is the formulas

OpenStudy (amistre64):

dx = (1--2)/n = (1+2)/n = 3/n

OpenStudy (anonymous):

k i got that far

OpenStudy (anonymous):

a=-2 b=1

OpenStudy (amistre64):

yep, them pictures is what I said lol

OpenStudy (amistre64):

the limit as n-> inf......

OpenStudy (anonymous):

so then I got lim as n goes to infinity of the summation [(-2+3i/n)^3-3(-2+3i/n)](3/n)

OpenStudy (amistre64):

need to work out something like this right?

OpenStudy (anonymous):

yea but that one is less complicated haha

OpenStudy (amistre64):

lol .... use the Esummation for the cubic function

OpenStudy (anonymous):

the answer we are going for is 3/4 for reference

OpenStudy (anonymous):

is my first step correct?

OpenStudy (amistre64):

we can split the addition up into to parts just like regualr integration, because regular integration is the result of this reimann sum stuff, so do x^3 and x sepertaely

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

[E] x^3 dx - [E] 3x dx the 3 can move out [E] x^3 (3/n) - 3 [E] x (3/n)

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\sum_{i=1}^{n}[(-2+3i/n)^3-3(-2+3i/n)](3/n)\]

OpenStudy (amistre64):

im lost on what that (-2+3i/n) stuff means, help me out :)

OpenStudy (anonymous):

its part of the reimann sum

OpenStudy (anonymous):

the formulas we have to use, that is the n^3

OpenStudy (anonymous):

i^3 my mistake

OpenStudy (amistre64):

\[\sum_{} x^{3}(3/n) -3 \sum_{}x(3/n)\] is whats in my head

OpenStudy (anonymous):

well the integral is \[\int\limits_{-2}^{1}(x^3-3x)dx\]

OpenStudy (amistre64):

yep, and the reiman is proof that we can split that into 2 parts right? \[\int\limits_{} x^3 dx -3\int\limits_{}x dx\]

OpenStudy (anonymous):

yes that works

OpenStudy (amistre64):

∑ x^3 (3/n)−3 ∑ x (3/n) is its equivalent, tell me if im wrong... but whats our next step after this?

OpenStudy (amistre64):

lets work each one on its own...ok

OpenStudy (anonymous):

plug in x^3 and x

OpenStudy (anonymous):

and thats what i showed above

OpenStudy (anonymous):

as one complete summation multiplied by 3/n

OpenStudy (amistre64):

you say "plug in" but i have no idea how that happens, im pretty much an idiot when it comes to the limit stuff, show me step by step please how we "plug in" to get: lim n→∞ ∑ i=1 n [(−2+3i/n)^3 −3(−2+3i/n)](3/n)

OpenStudy (anonymous):

k let me show you

OpenStudy (anonymous):

hopefully this helps

OpenStudy (anonymous):

i need to break it even further for you

OpenStudy (amistre64):

lol.... well, I got that part :) How do we move from that to the "formula for" [E] i^3 and [E] i I am not familiar with those formulas, yet :)

OpenStudy (anonymous):

its just pulling things out and taking individual summations, should i show you where i am so far?

OpenStudy (amistre64):

yes please ∑ i^3 = (1/4)n4 + (1/2)n3 + (1/4)n2 is what I found online

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\sum_{i=1}^{n}3/n[-8+12i/n+12i/n-18i^2/n^2+12i/n-18i^2/n^2-18i^2/n^2-27i^3/n^3-3(-2+3i/n)\]

OpenStudy (anonymous):

wow i guess my thing was too long, thats only half of it

OpenStudy (anonymous):

im so confused lol

OpenStudy (amistre64):

:) heres what I found that helps me out

OpenStudy (anonymous):

ok that works, lets start over then from 3/n being delta x

OpenStudy (amistre64):

since 3/n = delta x, we simply multiply that to x^3? to get: 3(x^3)/n ?? is that right?

OpenStudy (anonymous):

hmm sure, keep going because i wont know where you are going till you go further

OpenStudy (amistre64):

\[3\sum_{n=1}^{\infty} (x^3)/n\] Does this look right?

OpenStudy (anonymous):

no but keep going

OpenStudy (amistre64):

lol ....if it dont look right, then why should I keep going :)

OpenStudy (anonymous):

im trying to make sense of where you are going

OpenStudy (amistre64):

....... you gotta drive some of the way, im lost here for the moment

OpenStudy (anonymous):

oh k , im trying to think...problem is i cant find the equation that would help you

OpenStudy (anonymous):

let me see...

OpenStudy (anonymous):

where are you lost?

OpenStudy (anonymous):

how i formulated the entire sum?

OpenStudy (anonymous):

A=lim_{N→∞}Δx∑_{j=1}^{N}f(a+jΔx)

OpenStudy (anonymous):

does this help

OpenStudy (anonymous):

i=j

OpenStudy (amistre64):

yup.... im still at the split and trying to apply the formula for [E] x^3 to the left part. the top part of the [E] is gonna be (b-a) right? which we determined to be 3.... but its not sinking in....

OpenStudy (anonymous):

heres a better look

OpenStudy (amistre64):

ok....so delta x is a constant that gets pulled out to the left ... am I seeing that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you see how the i is appearing now. In the equation, it names it j, but the common known term is i

OpenStudy (amistre64):

and the "i" is simply each iteration of the sums....right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the integral should =3/4 after the long process, im trying to get through all the steps

OpenStudy (anonymous):

thats the problem

OpenStudy (amistre64):

yeah....when i read over that reaimann stuff I gloosed at it and figured, its true, so why do it the long way :)

OpenStudy (anonymous):

lol cause my professor doesnt want to make life easy

OpenStudy (amistre64):

f(a + i /x\) the i is each iterations from 1 to infinity. is the a the same as in our b-a interval?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

limn→∞∑i=1n[(−2+3i/n)3−3(−2+3i/n)](3/n)

OpenStudy (amistre64):

f(-2 + 3i/n) is a standard interation then, is that correct?

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

dont forget the cubed

OpenStudy (amistre64):

is the "n" from 3/n is the one we are limiting to infinity right?

OpenStudy (anonymous):

yes perfect

OpenStudy (amistre64):

f^3(x)...yeah yeah

OpenStudy (anonymous):

hey if you dont wanna type the dumb sigma and lim n thing, just say lim and sigma, ill know what you are saying or use the equation feature on here

OpenStudy (anonymous):

dumb summation sign i mean*

OpenStudy (amistre64):

n[(−2+3i/n)3−3(−2+3i/n)](3/n) this is you combining the terms together right??

OpenStudy (anonymous):

the n in front shouldnt be there that was part of the format for the summation

OpenStudy (amistre64):

ok...... (sigma) (-2 +3i/n) (3/n) is the right term then correct?

OpenStudy (anonymous):

summation but yes that is correct

OpenStudy (anonymous):

isnt it -3(-2 +3i/n) (3/n)

OpenStudy (anonymous):

you forgot the -3x

OpenStudy (amistre64):

\[\sum_{} (-2+3i/n) (3/n)\] would be the \[\sum_{} x \] formula part right? the -3 part of the -3x doesnts need to be included in this process if i recall correctly

OpenStudy (anonymous):

im pretty sure you do because its part of the integral function